OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 39 ?

sam (.sam.):

a6=a1+(n-1)d=39 (4)+(5)d=39

sam (.sam.):

d=7

sam (.sam.):

then , 32th term, a32=a1+(n-1)d 4+(31)(7)=?

OpenStudy (anonymous):

221

OpenStudy (precal):

The important thing to remember is that if you are told "arithmetic sequence or series", then you need to use thos particular formulas

sam (.sam.):

@dmartinez130 yes